OpenStudy (anonymous):

Simplify by combining like terms. x+x= 2+3+x= 2x+x+4+1= Please explain.........

jimthompson5910 (jim_thompson5910):

x+x = 2x since 1+1 = 2 (you basically add the coefficients)

jimthompson5910 (jim_thompson5910):

2+3 +x = 5+x The like terms are 2 and 3. I'll let you do the last one

OpenStudy (anonymous):

7+x?

jimthompson5910 (jim_thompson5910):

2x and x are like terms 4 and 1 are like terms

jimthompson5910 (jim_thompson5910):

2x+x = ??? 4+1 = ???